# Encourage or Inhibit Monosemanticity? Revisit Monosemanticity from a Feature Decorrelation Perspective

### v1

@@ Abstract
To better interpret the intrinsic mechanism of large language models (LLMs), recent studies focus on monosemanticity on its basic units. A monosemantic neuron is dedicated to a single and specific concept, which forms a one-to-one correlation between neurons and concepts. Despite extensive research in monosemanticity probing, it remains unclear whether monosemanticity is beneficial or harmful to model capacity. To explore this question, we revisit monosemanticity from the feature decorrelation perspective and advocate for its encouragement. We experimentally observe that the current conclusion by Wang et al. (2024), which suggests that decreasing monosemanticity enhances model performance, does not hold when the model changes. Instead, we demonstrate that monosemanticity consistently exhibits a positive correlation with model capacity, in the preference alignment process. Consequently, we apply feature correlation as a proxy for monosemanticity and incorporate a feature decorrelation regularizer into the dynamic preference optimization process. The experiments show that our method not only enhances representation diversity and activation sparsity but also improves preference alignment performance.
@@ 1 Introduction
Recent years have witnessed significant breakthroughs made by large language models (LLMs), and these LLMs demonstrate impressive performance across a wide range of NLP tasks (Rafailov et al., 2023b; Touvron et al., 2023; OpenAI, 2024). Meanwhile, understanding how they iteratively develop and refine suitable representations from inputs remains opaque (Zhou et al., 2024; Lee et al., 2024; He et al., 2024). Mechanistic interpretability is to understand neural networks by breaking them into components that are more easily understood than the entire network (Zhou et al., 2024; Lee et al., 2024; He et al., 2024). However, the neuron, the most basic computational unit of the neural network, is not a natural unit for human understanding. This is because many neurons are polysemantic, responding to mixtures of seemingly unrelated inputs (Bills et al., 2023; Gurnee et al., 2023; He et al., 2024).
@@
Towards fundamental interpretability, very recent works study the monosemantic neurons: those form a one-to-one correlation with their related input features  (Templeton et al., 2024; Bricken et al., 2023; Gurnee et al., 2023). Researchers in OpenAI have applied the sparse autoencoder (Cunningham et al., 2023) with dictionary learning to identify the monosemanticity at a large scale. Given the computational cost in training sparse autoencoder and the human labor required for generating interpretations, their detailed interpretability is specifically focused on 4096 features (Bricken et al., 2023). Furthermore, the study by Gurnee et al. (2023); Wang et al. (2024) proposes efficient monosemanticity proxies, offering a pathway for the exploration of this model property. Despite existing success, the relationship between monosemanticity and LLM’s capacity (such as robustness and alignment), remains a subject of ongoing debate. It raises an open question: Should monosemanticity be encouraged or inhibited for LLM’s alignment?
@@
To tackle the challenges mentioned above, in this paper, we revisit monosemanticity from the perspective of feature decorrelation and show a positive correlation between monosemanticity and within-model capacity. Consequently, we demonstrate this experimentally and propose a decorrelation regularization approach to enhance monosemanticity. Specifically, the main contributions of this paper are summarized as follows:
@@
  • (i) We review the recent studies in monosemanticity probing and identify the gap between existing qualitative analysis and quantitative optimization objectives.
@@
  • (ii) Our experiments show that while the relation between monosemanticity and cross-model capacity is unreliable, it is reliable within a single model, i.e., applying Direct Preference optimization (Rafailov et al., 2023a) (DPO) consistently improves monosemanticity (in Figure 2).
@@
  • (iii) We connect feature decorrelation with monosemanticity via the activation sparsity, thereby employing a decorrelation regularization to enhance monosemanticity. The co-occurrence of enhancement in activation sparsity and monosemanticity implies the validity of this connection.
@@
  • (iv) We apply this regularization to DPO, achieving efficient and robust preference alignment along with increased representation diversity and monosemanticity, further evidenced by a larger reward margin.
@@ 2 Monosemanticity Definition
[…]
Despite the utility of polysemantic neurons, to better interpret neural networks, more studies are focusing on the monosemanticity probing. In Contrast to the one-to-many mapping of polysemantic neurons, monosemantic neurons form a one-to-one correlation with their related input features. In addition to the interpretability of an individual neuron, monosemanticity also offers a novel perspective on disentanglement, sparsity, and scalability (Bricken et al., 2023; Gurnee et al., 2023; Wang et al., 2024).
Sparse AutoEncoder for semantics decompose.
@@
Identify monosemanticity at scale.
After decomposing the activation, we need to interpret each $\bm{f}_{i}$ and link it to a feature from a predefined disjoint feature set $\{A_{i}\}$. This feature set separates the input $\rm{\textbf{x}}\in X$ into $m$ features as follows:
@@
A neuron $\bm{z}$ is considered monosemantic if it is only activated by inputs that share a specific feature $A_{j}$ (Wang et al., 2024), that is:
$\forall_{\mathbf{x}}\text{act}(\bm{z},\mathbf{x})=1,\mathbf{x}\in A_{j}.$
However, these methods face two challenges that hinder the measurement of model-level monosemanticity and raise questions about monosemanticity optimization: (i) Each interpretation requires manual human analysis, prompting an advanced LLM with all the input text samples that activate $\bm{f}_{i}$ for interpretation and activation prediction (Bricken et al., 2023; Bills et al., 2023), making it difficult to be conducted at a large scale (Templeton et al., 2024). (ii) It is unclear if there is a ground truth or optimization objective for monosemanticity optimization. All the optimizations are only proposed for the sparse AutoEncoder training (Gao et al., 2024).
@@ 3 Monosemanticity Proxy
[…]
@@ 3.1 Unreliable evidence from superposition decomposition
The evidence inspiring their proposed inhibition hypothesis is presented in Figure 2 (c) of  Gurnee et al. (2023), which shows the layerwise product (defined in Eq. (4)) across multiple Pythia models (Biderman et al., 2023). The monosemanticity degree in Pythia-410M is higher than that in Pythia-6.9B. However, the monosemanticity in Pythia-1.4B is higher than that in Pythia-1B. So, there is no clear correlation between monosemanticity degree and model size. To further investigate this correlation, we applied this metric to GPT2-variants and the results are shown in Figure 2. When comparing GPT-2 variants with different parameter sizes, GPT-2 xl and GPT-2 large demonstrate greater overall monosemanticity than GPT-2 medium, with parameter sizes of 1.5B, 774M, and 355M, respectively, although the difference between GPT-2 large and GPT-2 xl is not significant. Therefore, we argue that there is no clear relation between the monosemanticity degree and model size. In fact, comparing different models may not be reliable due to numerous discrepancies, such as training data and training strategies.
Figure 2: Measured monosemanticity using product of the input weight norm $W_{fc}$ and bias $b_{fc}$ in the GPT2-based models. DPO consistently increases the monosemanticity degree over three GPT-2 variants (dashed line).
@@ 3.2 Understanding monosemanticity via decorrelation perspective
Based on the unreliable cross-model evidence in superposition decomposition, we now discuss the monosemanticity within models using activation sparsity.
As there are no bias terms in the MLPs weights of Llama-family models, the proxy in  Eq. (4) can not be applied. Wang et al. (2024) proposed an online model agnostic metric in Eq. (5) for assessing monosemanticity, which provides a theoretical guarantee given the computation limitations of existing probing methods.
$\phi(\bm{z}^{(m+1)})=\frac{(\bm{z}^{(m+1)}-\bar{\bm{z}})^{2}}{S^{2}},$ (5)
where $\bar{\bm{z}}$ is the average of historical samples across $m$ different inputs $\{\bm{z}^{[j]}\}_{j=1}^{(m)}$, and $S^{2}=\frac{\sum_{j=1}^{m}(\bm{z}^{[j]}-\bar{\bm{z}})^{2}}{m-1}$ is a constant for any $\bm{z}$. Therefore, this metric can be simplified to the activation variance across different input text samples.
@@
Highly correlated intermediate representations in language models.
According to Eq. (5), the more current activation deviates from the average activations obtained by other samples, the higher the monosemanticity. That is, we can take the feature decorrelation as a proxy of monosemanticity. In literature, highly correlated (less distinct) representations are a common issue observed in Transformer-based models due to the convex hull in self-attention (Yan et al., 2022; Dong et al., 2023).
@@
Recall the definition of superposition activation, where activations are linear combinations of multiple neurons, implying a high correlation among them. These non-orthogonal representations can also cause loss-increasing “interference” (Elhage et al., 2022). Recent works in toy models demonstrate that this tension manifests in a spectrum of representations: optimal capacity allocation tends to monosemantically represent the most important features, while polysemantically representing less important features (Scherlis et al., 2022).
3.3 Positive correlation between DPO and monosemanticity.
@@
Based on the two monosemanticity proxies, we now explore how the monosemanticity trends during the preference alignment process within the current language model.
@@
DPO enhances the monosemanticity degree based on superposition decomposition.
[…]
DPO increases the feature decorrelation.
@@
To study the characteristics of models without a bias term, we use feature decorrelation metric, i.e., (1-cosine similarity between activations from different inputs). Specifically, we train Llama on three datasets (details in § 5) via DPO and derive the MLP activation from 1000 randomly sampled input text from the corresponding dataset. We observe a clear promotion in dashed lines (DPO) in Figure 3.
@@
Figure 3: Feature decorrelation measurement over activations from Llama-2-7b-hf model. The activations are from the base model (inference on a specific dataset) and DPO (post-training on the same dataset). A well-trained DPO significantly increases the feature decorrelation, i.e., the proxy for monosemanticity.
@@ 4.1 Learn decorrelated activations
[…]
Figure 4: Feature decorrelation measurement across different layers in Llama2-7b-hf during the preference optimization process. The number in the name of each curve represents the training step. Both DPO and DecPO greatly increase the feature decorrelation over Base(0-step) very quickly, followed by a pronounced overfitting widely studied in the literature. DecPO achieves significantly higher decorrelation, especially in the late training stage, thereby reducing the speed of overfitting.
@@ 5 Monosemanticity Contributes to Preference Optimization
The previous section has provided evidence that a decorrelation regularizer can enhance monosemanticity. Now, we continue to validate our hypothesis, monosemanticity should be encouraged, by evaluating whether DecPO will boost alignment performance. Although decorrelated representations have been widely discussed in both computation vision and language processing (Hua et al., 2021; Yan et al., 2023), limited research has examined this issue within existing preference optimization algorithms, such as DPO (Rafailov et al., 2023a) and Proximal Policy Optimization (PPO) (Schulman et al., 2017).

### v2

@@ Abstract
To better interpret the intrinsic mechanism of large language models (LLMs), recent studies focus on monosemanticity on its basic units. A monosemantic neuron is dedicated to a single and specific concept, which forms a one-to-one correlation between neurons and concepts. Despite extensive research in monosemanticity probing, it remains unclear whether monosemanticity is beneficial or harmful to model capacity. To explore this question, we revisit monosemanticity from the feature decorrelation perspective and advocate for its encouragement. We experimentally observe that the current conclusion by Wang et al. (2024), which suggests that decreasing monosemanticity enhances model performance, does not hold when the model changes. Instead, we demonstrate that monosemanticity consistently exhibits a positive correlation with model capacity, in the preference alignment process. Consequently, we apply feature correlation as a proxy for monosemanticity and incorporate a feature decorrelation regularizer into the dynamic preference optimization process. The experiments show that our method not only enhances representation diversity but also improves preference alignment performance  [^1] The code is released at https://github.com/hanqi-qi/Revisit_monosemanticity..
@@ 1 Introduction
Recent years have witnessed significant breakthroughs made by large language models (LLMs), which demonstrate impressive performance across a wide range of NLP tasks (Rafailov et al., 2023; Touvron et al., 2023; OpenAI, 2024). Meanwhile, understanding how they iteratively develop and refine suitable representations from inputs remains opaque (Zhou et al., 2024; Lee et al., 2024; He et al., 2024). Mechanistic interpretability is to understand neural networks by breaking them into components that are more easily understood than the entire network (Zhou et al., 2024; Lee et al., 2024; He et al., 2024). However, the neuron, the most basic computational unit of the neural network, is not a natural unit for human understanding. This is because many neurons are polysemantic, responding to mixtures of seemingly unrelated inputs (Bills et al., 2023; Gurnee et al., 2023; He et al., 2024).
@@
Towards fundamental interpretability, very recent works study the monosemantic neurons: those form a one-to-one correlation with their related input features  (Templeton et al., 2024; Bricken et al., 2023; Gurnee et al., 2023). Researchers in OpenAI have applied the sparse autoencoder (Cunningham et al., 2023) with dictionary learning to identify the monosemanticity at a large scale. Given the computational cost in training sparse autoencoder and the human labor required for generating interpretations, their detailed interpretability is specifically focused on 4,096 features (Bricken et al., 2023). Furthermore, the studies by Gurnee et al. (2023) and Wang et al. (2024) proposed efficient monosemanticity proxies, offering a pathway for the exploration of this model property. Despite success, the relationship between monosemanticity and LLM’s capacity (such as robustness and alignment), remains a subject of ongoing debate. It raises an open question: Should monosemanticity be encouraged or inhibited for LLM’s alignment?
@@
To tackle the aforementioned challenges, in this paper, we revisit monosemanticity from the perspective of feature decorrelation and show a positive correlation between monosemanticity and within-model capacity. Consequently, we demonstrate this experimentally and propose a decorrelation regularization approach to enhance monosemanticity. Specifically, the main contributions of this paper are summarized as follows:
@@
  • (i) We have reviewed recent studies in monosemanticity probing and identified the gap between current qualitative analysis and quantitative optimization objectives.
@@
  • (ii) Our experiments show that while the relationship between monosemanticity and cross-model capacity is inconsistent, it is reliable within a single model, specifically applying Direct Preference optimization (Rafailov et al., 2023) (DPO) consistently improves monosemanticity, as shown in Figure 2.
@@
  • (iii) We establish a link between feature decorrelation and monosemanticity through activation sparsity, employing decorrelation regularization to enhance monosemanticity. The concurrent enhancement in activation sparsity and monosemanticity supports the validity of this connection.
@@
  • (iv) We implement this regularization with DPO, achieving efficient and robust preference alignment alongside increased representation diversity and monosemanticity, as further evidenced by a larger reward margin.
@@ 2 Monosemanticity Definition
To avoid confusion caused by terminology, we first clarify the definitions of the terms concept, feature, and neuron in this context.
  • Concept in our paper refers to an interpretable property of the input that would be recognizable to most humans.
  • Neuron refers to a node in a neural network, associated with model weights.
  • Features are the representation or activation to refer to the model intermediate vector/outputs.
The challenge of explaining neurons lies in the fact that many of them are polysemantic: they respond to mixtures of distinct concepts, i.e, $n$ concepts in $d<n$ dimensions. It naturally arises in the neural network (NN) training process as more high-level intermediate features are aggregated by combining the neurons of the NN. Despite the utility of polysemantic neurons, to better interpret neural networks, more studies are focusing on the monosemanticity probing. In Contrast to the one-to-many mapping of polysemantic neurons, monosemantic neurons form a one-to-one correlation with their related input features. In addition to the interpretability of an individual neuron, monosemanticity also offers a novel perspective on disentanglement, sparsity, and scalability (Bricken et al., 2023; Gurnee et al., 2023; Wang et al., 2024).
Sparse AutoEncoder for semantics decomposition.
@@
Identifying monosemanticity at scale.
After decomposing the activation, we need to interpret each $\bm{f}_{i}$ and link it to a concept from a predefined disjoint concept set $\{A_{i}\}$. This concept set divides all input samples $X$ into $m$ concepts, where each input $x$ is considered to represent a single concept (e.g., past tense):
@@
A neuron $\bm{z}$ is considered monosemantic if it is only activated by inputs that share a specific concept $A_{j}$ (Wang et al., 2024), that is:
$\forall_{\mathbf{x}}\text{activation}(\bm{z},\mathbf{x})=1,\mathbf{x}\in A_{j}.$
However, these methods face two challenges that hinder the measurement of model-level monosemanticity and raise questions about monosemanticity optimization: (i) Each interpretation requires manual human analysis, involving prompting an advanced LLM with all the input text samples that activate $\bm{f}_{i}$ for interpretation and activation prediction (Bricken et al., 2023; Bills et al., 2023), making it difficult to conduct at a large scale (Templeton et al., 2024). (ii) It is unclear whether there is a ground truth or optimization objective for monosemanticity. Currently, optimizations are only proposed within the context of sparse AutoEncoder training (Gao et al., 2024).
@@ 3 Monosemanticity Proxy
Llama-family (Touvron et al., 2023) models introduce an extra $W_{\text{gate}}$ and omit all the bias terms in the weight matrix: [^2] We use the same symbol as the Llama source code for weight matrices.
@@ 3.1 Unreliable evidence from superposition decomposition
The evidence inspiring their proposed inhibition hypothesis is presented in Figure 2 (c) of  Gurnee et al. (2023), which shows the layerwise product (defined in Eq. (4)) across multiple Pythia models (Biderman et al., 2023). The monosemanticity degree in Pythia-410M is higher than that in Pythia-6.9B. However, the monosemanticity in Pythia-1B is lower than that in Pythia-1.4B. So, there is no clear correlation between monosemanticity degree and model size. To further investigate this correlation, we applied this metric to GPT2-variants and show the results in Figure 2. When comparing GPT-2 variants with different parameter sizes, GPT-2 xl (1.5B) and GPT-2 large (774M) demonstrate greater overall monosemanticity than GPT-2 medium (355M), although the monosemanticity of GPT-neo-2.7B is lower than that of the aforementioned GPT-2 variants. Therefore, we argue that there is no clear relationship between the monosemanticity degree and the model size. In fact, comparing different models may not be reliable due to numerous discrepancies, such as training data and training strategies.
Figure 2: Measured monosemanticity using product of the input weight norm $W_{fc}$ and bias $b_{fc}$ in the GPT2-based models. There is no consistent correlation between monosemanticity and model sizes.
@@ 3.2 Understanding monosemanticity via decorrelation perspective
Based on the inconsistent cross-model evidence in superposition decomposition, we now discuss the monosemanticity within models using feature decorrelation via a theoretical justification.
Theoretical justification of the relationship between monosemanticity and decorrelation.
In the seminal work (Elhage et al., 2022), the researchers in Anthropic identified superposition as a critical source of polysemanticity (i.e., the opposite of monosemanticity). Superposition refers to the phenomenon where the models encode more features than the number of neurons, such that it is impossible for different neurons not to interface (non-orthogonal) with each other. This motivates Elhage et al. (2022) to use model weight (associated with the neuron) correlation as a measure of superposition. In particular, with a linear toy model, they use the following correlation as measurement for superposition:
$\sum_{j\neq i}\left(\bm{W}_{i}\cdot\bm{W}_{j}\right)^{2},$
where $\bm{W}_{i}$ and $\bm{W}_{j}$ are two different model weight vectors, to essentially measure the off-diagonal terms of the correlation matrix $\bm{W}^{\top}\bm{W}$. If the neurons are monosemantic (uncorrelated), then $\bm{W}^{\top}\bm{W}$ would be a diagonal matrix $\bm{D}$.
Following this definition, in large language models, we measure the correlation between the feature/activation $z$ to measure superposition at different layers. It is easy to see that the activation correlation is equivalent to the Anthropic’s measure under linear models and independent features (considered in their paper). Let $\bm{Z}=\bm{WX}$ be the activation of the linear model, where $\bm{X}$ is the input, $\bm{W}$ is the weight matrix. if we have $\bm{W}^{\top}\bm{W}=\bm{D}$ and $\bm{X}^{\top}\bm{X}=\bm{I}$, we will have:
$\bm{Z}^{\top}\bm{Z}=\bm{X}^{\top}\bm{W}^{\top}\bm{WX}=\bm{D}.$
Thus, $\bm{Z}^{\top}\bm{Z}$ is a diagonal matrix when the neurons are uncorrelated, i.e., monosemantic. Driven by this connection, we develop the feature/activation decorrelation loss between normalized activations (whose diagonal terms are 1) as our proxy and regularisation loss. Therefore, there is indeed a close connection between our feature decorrelation loss and monosemanticity phenomenon.
@@
Highly correlated intermediate representations are commonly observed in language models.
@@
In literature, highly correlated (less distinct) representations are a common issue observed in Transformer-based models due to the convex hull in self-attention (Yan et al., 2022; Dong et al., 2023). Recall the definition of superposition activation, where activations are linear combinations of multiple neurons, implying a high correlation among them. These non-orthogonal representations can also cause loss-increasing “interference” (Gurnee et al., 2023). Recent works in toy models demonstrate that this tension manifests in a spectrum of representations: optimal capacity allocation tends to monosemantically represent the most important features, while polysemantically representing less important features (Scherlis et al., 2022).
3.3 Positive correlation between DPO and feature decorrelation.
@@
Based on the monosemanticity proxy, i.e., decorrelation, we investigate the trends in monosemanticity during the preference alignment process within the current language models.
@@
DPO enhances the monosemanticity degree based on superposition decomposition, especially in the earlier layers.
We implemented DPO on the three variants of GPT-2 and measured the monosemanticity degree using the product method. [^3] As Llama-family models do not have a bias term, the product method cannot be applied to them. We selected the top 100 dimensions of $W_{in}$ because most parameters exhibit minimum changes after DPO, consistent with observations in Lee et al. (2024). The results after DPO are in shown in Figure 4. DPO training indeed improves monosemanticity in earlier model layers, the this effect is consistent across different GPT-2 models. The decline in monosemanticity observed in later layers can be attributed to the increased complexity and polysemantic nature of information nearer to the prediction layer, which is necessary for handling diverse tasks. This finding is consistent with that of  (Lee et al., 2024). They identified several MLP dimensions as toxicity vectors in $\text{GPT}_{\text{DPO}}$, and after subtracting these vectors, they observed a significant decrease in toxicity of the generated text. This change was much less evident in stanard GPT models. This suggests that DPO training makes certain dimensions more responsive to specific features, a characteristic that reflects monosemanticity (Further evidence is provided in §5, Table 1).
DPO increases feature decorrelation.
@@
To study the characteristics of models without a bias term, we use the feature decorrelation metric, defined as (1--cosine similarity between activations from different inputs), as a proxy for monosemanticity. Specifically, we train Llama on three datasets (details in § 5) using DPO and extract the MLP activations from 1,000 randomly sampled input texts from each respective dataset. We observe a clear enhancement in the dashed lines (representing DPO) in Figure 4. The trends in feature decorrelation is similar to that seen in Figure 4, which empirically validates the use of feature decorrelation as a proxy for monosemanticity.
Figure 3: Relative changes of normalized median $(||w_{in}||_{2}b_{in})$, a proxy for monosemanticity, across different GPT2 models after DPO training.
@@
Figure 4: Feature decorrelation measurement of activations from the Llama-2-7b-hf model. The activations are derived from both the base model (inference on a specific dataset) and DPO (post-training on the same dataset). A well-trained DPO significantly increases feature decorrelation, i.e., the proxy for monosemanticity. The drop in later layers has also been observed in Yan et al. (2022), attributed to their proximity to the supervision signal.
@@ 4.1 Learn decorrelated activations
We apply DecPO to Llama2-7b-hf  [^4] https://huggingface.co/meta-llama/Llama-2-7b-hf on the Toxicity dataset (Lee et al., 2024). The results of representation decorrelation at various training stages are shown in Figure 5. We observe a significant and rapid increase in feature decorrelation for both DPO and DecPO compared to the Base model, followed by a decrease, implying an overfitting issue widely observed in previous studies (Deng et al., 2023; Azar et al., 2024; Pal et al., 2024). Additionally, DecPO significantly reduces the overfitting speed, demonstrated by the smaller gaps between different dashed lines compared to the solid ones. The enhancement from DecPO is more pronounced in the late stage of training.
Figure 5: Feature decorrelation measurement across different layers in Llama2-7b-hf during the preference optimization process. The number in the name of each curve represents the training step. Both DPO and DecPO greatly increase the feature decorrelation over Base(0-step) very quickly, followed by a pronounced overfitting widely studied in the literature. DecPO achieves higher decorrelation, especially in the late training stage, thereby reducing the speed of overfitting.
@@ 5 Monosemanticity Contributes to Preference Optimization
The previous section has provided evidence that a decorrelation regularizer can enhance monosemanticity. Now, we continue to validate our hypothesis, monosemanticity should be encouraged, by evaluating whether DecPO will boost alignment performance. Although decorrelated representations have been widely discussed in both computation vision and language processing (Hua et al., 2021; Yan et al., 2023), limited research has examined this issue within existing preference optimization algorithms, such as DPO (Rafailov et al., 2023) and Proximal Policy Optimization (PPO) (Schulman et al., 2017).